Reports 1-1 of 1 Clear search Modify search
MIF (Noise Budget)
takafumi.ushiba - 21:52 Friday 24 February 2023 (24152) Print this report
TF measurement of PRCL2DARM and MICHDARM for updating noise budget.

Since current noise budget from MICH and PRCL are calculated TFs before engaging feed forward, I measured the TF from MICH2DARM (fig1) and PRCL2DARM (fig2) with FFs.
I will update the noise budget soon.

Images attached to this report
Comments to this report:
takafumi.ushiba - 13:52 Saturday 25 February 2023 (24156) Print this report

I updated the noise budget in medm shown at control room.
Currently, Type-A DAC noise is well below the sensitivity and MICH and PRCL noise are close but slighty below the sensitivity.
See here for the updated noise budget.

Search Help
×

Warning

×